I need 10 different expressions that end up equaling a different number, the answer must be between 1-20 and use the order of operations. You may only use five 5's and no other numbers.
1
Expert's answer
2011-11-04T09:46:03-0400
5+5*5-5-5=20
5+5-(5+5)/5=7
5+5+5-5+5=15
(5+5*5)/(5+5)=3
(5+5*5)/5+5=11
5*5/5+5-5=5
5/5+5*5/5=6
5/5+5/5+5=7
5*(5-5)+5/5=1
5/5*(5+5)/5=2